# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v2

@@ Abstract
Gradient-based optimization is the workhorse of deep learning, offering efficient and scalable training via backpropagation. However, exposing gradients during training can leak sensitive information about the underlying data, raising privacy and security concerns such as susceptibility to data poisoning attacks. In contrast, black box optimization methods, which treat the model as an opaque function, relying solely on function evaluations to guide optimization, offer a promising alternative in scenarios where data access is restricted, adversarial risks are high, or overfitting is a concern. This paper introduces BBoxER, an evolutionary black-box method for LLM post-training that induces an information bottleneck via implicit compression of the training data. Leveraging the tractability of information flow, we provide non-vacuous generalization bounds and strong theoretical guarantees for differential privacy, robustness to data poisoning attacks, and extraction attacks. In experiments with LLMs, we demonstrate empirically that black-box optimization methods—despite the scalability and computational challenges inherent to black-box approaches—are able to learn, showing how a few iterations of BBoxER improve performance, generalize well on a benchmark of reasoning datasets, and are robust to membership inference attacks. This positions BBoxER as an attractive add-on on top of gradient-based optimization, offering suitability for deployment in restricted or privacy-sensitive environments while also providing non-vacuous generalization guarantees.
@@ 1 Introduction
We introduce BBoxER (Black-Box Evolutionary Retrofitting), a comparison-based black-box retrofitting method [30] applicable after pretraining, fine-tuning, or reinforcement learning loops such as GRPO [31]. Positioned in the outermost layer of the pipeline, BBoxER requires no gradient access and integrates seamlessly with existing black-box libraries and algorithms [32]. The resulting sequence of queries and updates forms an implicit compression trace, enabling the use of compression-based generalization theory [33, 34], currently the only route to non-vacuous generalization bounds for LLMs [35]. It allows us to derive strong generalization and differential privacy guarantees that - unlike VC dimension [36] or Rademacher complexity [37] - do not depend on the number of modified parameters but on the complexity of the optimization trajectory. As a result, BBoxER ensures privacy by design, robustness to data poisoning, robustness to extraction attacks, and provable, non-vacuous generalization bounds to control overfitting. It is uniquely positioned to provide improvements from aggregate feedback of billions of users anonymously, without accessing individual data.
@@
$\bullet$ Strong theoretical guarantees. We derive non-vacuous generalization bounds (Corollary˜1) that scale linearly with dataset size, and establish privacy by design through provable robustness to poisoning (Equation˜7) and resistance to extraction attacks (Corollary˜2) (Secs.˜4, 4.2 and 4.3).
@@ 2 Background and Related Work
Privacy and Robustness. LLMs risk leaking private or confidential content [4]. Differential privacy (DP) techniques [44] like DP-SGD [45, 46] and PATE [47] are the standard paradigm for privacy preserving learning, offering formal guarantees but suffering from utility and scalability challenges, due to the necessary per-sample gradient clipping and adding noise that is inversely proportional to $\varepsilon$. Our method avoids the need for noisy gradient updates by enforcing a compression bottleneck through comparison-based optimization. For instance, in the specific case of ensuring the privacy of user input prompts and outputs in A/B testing, it achieves $(\varepsilon=0,\delta=0)$.
@@ 3 Methods: Retrofitting with BBO
Retrofitting [48, 49, 50] is routinely deployed in industry, in order to adapt old devices to new contexts or new data. In machine learning, the term has mainly been used in natural language processing since Faruqui et al. [51], who adapted the representation of the word vector to take semantic knowledge into account. It is performed after classical training, i.e., without retraining the entire network. Videau et al. [30] adapted retrofitting to the black-box optimization context. In the present paper, we instantiate retrofitting as BBoxER and consider comparison-based black-box optimization algorithms to obtain the explicit compression properties that lend themselves to the strong bounds proved in Sec.˜4.
@@ 3.1 The Retrofitting Framework: BBoxER
First, Algorithm $a$ is initialized and returns its initial internal state $I_{0}$ in ˜1. Then for $b$ iterations, Algorithm $a$, via $a.ask$ at ˜3, proposes a new sample $x_{i}$, which used to obtain a new model $m_{i}$ through modifiedmodified. ˜4 introduces the branching factor $k_{i}$, representing the finite number of possible outputs of $a.compare$, which we assume is upper bounded. At ˜5, $a.compare$ returns $choice_{i}\in\{1,\dots,k_{i}\}$, the outcome of comparing various previously proposed models $m_{1},\dots,m_{i}$ on the data $D$. For instance, the comparison may involve the current model and the best model obtained so far, which corresponds to $k_{i}=2$. The outcome of this comparison can be based on user-provided preference, such as in A/B testing, or on performance evaluations over $D$, where we compute a metric (e.g., accuracy) but only retain the binary comparison result (e.g., 0 or 1). In ˜7, internal state of $a$ is updated to account for the comparison result. Finally $a.recommend$ at ˜9 returns the modification $\widehat{x}$ used to obtain the final model.
It is important to note that the internal state $I_{b}$ of algorithm $a$ is entirely determined by $\omega,a,b,$ $choice_{1},\dots,choice_{b}$ and contains all the information needed to obtain $\widehat{x}$, and thus the final modified model. Moreover, the number of distinct possible outputs of BBoxER is upper bounded by the number of distinct possible values of $I_{b}$. We assume that the randomness of algorithm $a$ is fixed given the seed $\omega$; stochastic algorithms can be obtained by randomizing the seed $\omega$ in $a$.
@@ 4.1 Overfitting bounds
Most BBO algorithms used within BBoxER bound the average $k_{i}$ by $2$ or $3$. Interpreting $\log_{2}(k_{i})$ as a budget in bits means that to avoid overfitting we can afford a number of iterations $b$ proportional to $s$ divided by the bits used per step. Precisely, if we assume $\forall i,k_{i}=2$, then for an error gap $\epsilon$ and confidence $\delta$, we have from Corollary˜1 the following formula (numerical values can be found in Sec.˜E.7):
@@ 4.2 Privacy and poisoning
Differential privacy guarantee Let $D_{1}$ and $D_{2}$ be two datasets that differ by a single sample. We examine the effect of this difference on the FinalModel given by BBoxER, and we distinguish two cases:
$\bullet$ privacy of users prompts and outputs (like in A/B testing) We assume that the preference signals are provided directly by users, and that the datasets $D_{1}$ and $D_{2}$ can differ on the prompts or outputs. Since BBoxER relies solely on users preference signals to perform model comparisons, and does not access prompts or outputs, privacy is ensured by design and we have $(\varepsilon,\delta)=(0,0)$.
$\bullet$ privacy of the entire sample We assume that the $D_{1}$ and $D_{2}$ can differ on an entire sample, which impacts its prefence signal. This is the case when the preference signal needs to be computed from the sample before aggregating and passing it to the algorithm. In this scenario, we obtain $\varepsilon=0$, while $\delta\;\propto\;\frac{b}{\sqrt{s}}$, following the poisoning bound from Thm˜2 with $m=1$.
Thm˜2 yields two immediate consequences: First, unless $m=\Omega(\sqrt{s})$, the retrofitting framework is resilient to poisoning attacks, and second, privacy is guaranteed by design, covering prompts, outputs, and user satisfaction signals without additional mechanisms. These points are further discussed in App.˜F. It is important to note that, while other privacy preserving methods such as DP-SGD [46] enforce privacy by perturbing gradients and consequently suffer from utility degradation as $\varepsilon$ becomes small, BBoxER’s compression-based guarantees remain robust. These structural differences highlight a fundamental advantage of our approach.
@@ 5.1 Experiments
Experiment 4 (Robustness to data extraction). To assess the robustness of BBoxER against membership inference attacks (MIAs), we begin from the observation that most MIAs rely on the negative log-likelihood (NLL) of training samples as their primary signal to infer membership [63, 64, 65] - see also App.˜H for an exposition. Accordingly, we propose to measure the average of absolute differences in negative log-likelihood (NLL) between the base and the updated model for each training sample. We consider GSM8K as our training dataset, Llama-3.1-8B as our base model, and compute three metrics: Diff-full (NLL computed over the entire prompt), Diff-CoT-a (NLL computed over only the CoT and the answer), and the relative average exact match improvement on GSM8K-test from the base model. For a fair comparison, we evaluate BBoxER alongside standard fine-tuning and DP-AdamW, training all methods on GSM8K under different settings-varying budget, epochs, and parameter subsets (full details in App.˜H)—so that models are compared at similar utility. Our results, reported in Tab.˜2, demonstrate that at matched test accuracy, the NLL shifts under BBoxER are orders of magnitude smaller than under fine-tuning or DP-AdamW. This provides empirical confirmation of our theoretical claim in Sec.˜4.3: BBoxER is significantly more robust to extraction and membership inference attacks.
@@ 5.2 Discussion of experimental results
Computational cost. Our approach is memory-efficient, as it relies solely on LLM inference without the need to compute and store gradients, or large optimizer states. This allows us to fit a large batch size during training. The overhead introduced by the internal state of the BBO algorithm is negligible compared to the computational cost of LLM inference.
@@ Appendix C Theoretical Background: A Brief Overview of Statistical Learning Theory
C.1 Non Asymptotic Large Deviation Bounds from the 60s: Hoeffding, Bennet, Bernstein
@@
In case we have a prior on a bound $\sigma^{2}$ on the variance of a randomly drawn individual loss, we can use the bound of Bennet and Bernstein [104]:
@@ C.2.1 Bounds on the Generalization Loss based on Bonferroni
Thm˜1 extends this to more sophisticated cases, with $N(\omega,a,b)=\pi_{i}k_{i}(\omega,a,b)>>b$ (in many cases, exponential in $b$ e.g. $N(\omega,a,b)\leq 2^{b}$ and $k_{i}\leq 2$) instead of $N(\omega,a,b)=b$. In spite of being larger, this bound, exponential in $b$, is tighter than bounds based on capacity.
@@ D.1 Evolutionary Algorithms
CMA-ES [13, 14, 15] iteratively updates a multivariate Gaussian random variable, initialized at the identity and parametrized by a covariance matrix and its variance (the so-called step-size). Each update is based on the correlations observed in successful mutations. This allows CMA-ES to adapt its step-size and search directions (the covariance matrix) to the problem landscape. Despite its $o(d^{2})$ computational cost and memory usage, due the the handling of the full covariance matrix, it is widely regarded as a powerful and reliable optimizer.
@@ E.2 Generalization Bounds for Randomized BBoxER
In contrast to random search, where all candidate parameterizations are chosen independently of their evaluation, comparison-based algorithms build candidates iteratively based on performance. If $\lambda$ candidates are evaluated at each step and the best is retained, then the number of internal states after one iteration is $\lambda$; after two iterations, $\lambda^{2}$; and after $n=b/\lambda$ iterations, at most $\lambda^{n}$. The total risk is, therefore, bounded by $\lambda^{n}\cdot\delta_{1,\epsilon}$, where $n$ is the number of iterations with $k_{i}>1$ (with $k_{i}=\lambda$ in such iterations); see [108].
@@ E.7 Dependence on Dataset Size: Scaling and Numerical Application
If we prefer the Bennet&Bernstein inequality (based on an assumption on the standard deviation, see Sec.˜C.1 [102, Section 8.2]), we get
@@
This equation, for $s=8000$ examples (close to GSM8k), $\delta=\frac{1}{2}$ (we are considering the median case), $\epsilon=0.01$ (we want that precision) and $\sigma=0.06$ (the standard variation of the loss is 0.06, we assume that the initial model is already good), leads to the claim that a budget $b\leq 91$ cannot lead to overfitting. If we assume $\epsilon=0.04$ and $\sigma=0.3$, we get with Hoeffding a bound $34$ and with Bennet&Bernstein a bound $88$. With $\epsilon=0.06$, $\sigma=0.3$, Bennet&Bernstein leads to a budget $189$. The budget becomes $482$ if we use $\epsilon=0.1$, $\sigma=0.3$. Also, these bounds scale linearly with $s$ (the number of user preferences in A/B testings or dataset size). These numbers are low, but we see that our bounds are not vacuous even in realistic scenarios.
@@ F.1 Privacy by Design
Privacy of the entire sample.
(general case of applying BBoxER on a dataset $D$) We may also be interested in the privacy implications of modifying a single satisfaction response $r_{i}$ associated with a specific user, or in the case were we compute this preference signal from the data. While this is less sensitive than access to the user’s actual queries or responses, it still reflects user preference and may warrant protection. In this setting, we consider the case where $D_{1}$ and $D_{2}$ differ by exactly one such response, altered during a run.
Applying Equation˜7 with $m=1$, we get that $A(D_{1})$ and $A(D_{2})$ are the same with probability at least $1-3b/\sqrt{2s\pi}$ when we have $s$ data points and $b$ iterations. This corresponds to differential privacy with $\varepsilon=0$ and $\delta=3b/\sqrt{2s\pi}$.

### v3

# Tuning without Peeking: Provable Generalization Bounds and Robust LLM Post-Training

@@ Abstract
[…]
@@ 1 Introduction
We introduce BBoxER (Black-Box Evolutionary Retrofitting), a comparison-based black-box retrofitting method [30] applicable after pretraining, fine-tuning, or reinforcement learning loops such as GRPO [31]. Positioned in the outermost layer of the pipeline, BBoxER requires no gradient access and integrates seamlessly with existing black-box libraries and algorithms [32]. The resulting sequence of queries and updates forms an implicit compression trace, enabling the use of compression-based generalization theory [33, 34], currently the only route to non-vacuous generalization bounds for LLMs [35]. It allows us to derive strong generalization and privacy guarantees that - unlike VC dimension [36] or Rademacher complexity [37] - do not depend on the number of modified parameters but on the complexity of the optimization trajectory. As a result, BBoxER ensures privacy by design, robustness to data poisoning, robustness to extraction attacks, and provable, non-vacuous generalization bounds to control overfitting. It is uniquely positioned to provide improvements from aggregate feedback of billions of users anonymously, without accessing individual data.
@@
$\bullet$ Strong theoretical guarantees. We derive non-vacuous generalization bounds (Cor.˜1) that scale linearly with dataset size, and establish privacy by design through provable robustness to poisoning (Equation˜7) and resistance to extraction attacks (Cor.˜2) (Secs.˜4, 4.2 and 4.3).
@@ 2 Background and Related Work
Privacy and Robustness. LLMs risk leaking private or confidential content [4]. Differential privacy (DP) techniques [44] like DP-SGD [45, 46] and PATE [47] are the standard paradigm for privacy preserving learning, offering formal guarantees but suffering from utility and scalability challenges, due to the necessary per-sample gradient clipping and adding noise that is inversely proportional to $\varepsilon$. Our method avoids the need for noisy gradient updates by learning directly from user preference signal, in the specific case of A/B testing, ensuring perfect privacy of users input prompts and outputs. Unlike SFT or RLHF, BBoxER ensures that the model never accesses any raw example, only aggregated preference signals. This aligns with emerging concerns in LLM safety [48, 49] about leakage during alignment, rather than database style privacy.
@@ 3 Methods: Retrofitting with BBO
Retrofitting [50, 51, 52] is routinely deployed in industry, in order to adapt old devices to new contexts or new data. In machine learning, the term has mainly been used in natural language processing since Faruqui et al. [29], who adapted the representation of the word vector to take semantic knowledge into account. It is performed after classical training, i.e., without retraining the entire network. Videau et al. [30] adapted retrofitting to the black-box optimization context. In the present paper, we adapt retrofitting to LLM post-training and instantiate it as BBoxER and consider comparison-based black-box optimization algorithms to obtain the explicit compression properties that lend themselves to the strong bounds proved in Sec.˜4.
@@ 3.1 The Retrofitting Framework: BBoxER
First, Algorithm $a$ is initialized and returns its initial internal state $I_{0}$ in ˜1. Then for $b$ iterations, Algorithm $a$, via $a.ask$ at ˜3, proposes a new sample $x_{i}$, which used to obtain a new model $m_{i}$ through modifiedmodified, which can be any function that modifies the model weights using the sampled noise by Algorithm $a$. Examples of such modifiedmodified function are provided in App.˜B. ˜4 introduces the branching factor $k_{i}$, representing the finite number of possible outputs of $a.compare$, which we assume is upper bounded. At ˜5, $a.compare$ returns $choice_{i}\in\{1,\dots,k_{i}\}$, the outcome of comparing various previously proposed models $m_{1},\dots,m_{i}$ on the data $D$. For instance, the comparison may involve the current model and the best model obtained so far, which corresponds to $k_{i}=2$. The outcome of this comparison can be based on user-provided preference, such as in A/B testing, or on performance evaluations over $D$, where we compute a metric (e.g., accuracy) but only retain the binary comparison result (e.g., 0 or 1). In ˜7, internal state of $a$ is updated to account for the comparison result. Finally $a.recommend$ at ˜9 returns the modification $\widehat{x}$ used to obtain the final model. A step by step execution example of BBoxER can be found in App.˜K.
It is important to note that the internal state $I_{b}$ of algorithm $a$ is entirely determined by $\omega,a,b,\allowbreak choice_{1},\dots,choice_{b}$ and contains all the information needed to obtain $\widehat{x}$, and thus the final modified model. Moreover, the number of distinct possible outputs of BBoxER is upper bounded by the number of distinct possible values of $I_{b}$. We assume that the randomness of algorithm $a$ is fixed given the seed $\omega$; stochastic algorithms can be obtained by randomizing the seed $\omega$ in $a$.
@@ 4.1 Overfitting bounds
Most BBO algorithms used within BBoxER bound the average $k_{i}$ by $2$ or $3$. Interpreting $\log_{2}(k_{i})$ as a budget in bits means that to avoid overfitting we can afford a number of iterations $b$ proportional to $s$ divided by the bits used per step. Precisely, if we assume $\forall i,k_{i}=2$, then for an error gap $\epsilon$ and confidence $\delta$, we have from Cor.˜1 the following formula (numerical values can be found in Sec.˜E.7):
@@ 4.2 Privacy and poisoning
Privacy guarantee Let $D_{1}$ and $D_{2}$ be two datasets that differ by a single sample. We examine the effect of this difference on the FinalModel given by BBoxER, and we distinguish two cases:
$\bullet$ privacy of users prompts and outputs (like in A/B testing) We assume that the preference signals are provided directly by users, and that the datasets $D_{1}$ and $D_{2}$ can differ on the prompts or outputs. Since BBoxER relies solely on users preference signals to perform model comparisons, and does not access prompts or outputs, privacy is ensured by design and we have $(\varepsilon,\delta)=(0,0)$. In this scenario, the user preference signal isn’t considered private, but the input prompts and outputs are private and they can be completely different without impacting the output of BBoxER.
$\bullet$ algorithm stability We assume that the $D_{1}$ and $D_{2}$ can differ on an entire sample. This is the case when the preference signal needs to be computed from the data before aggregating and passing it to the algorithm or when we consider the user’s preference signal to be a private information in A/B testing. In the following we bound the influence of one sample. From Equation˜7 using $m=1$ (only 1 sample difference) we get that : $P\!\left(output_{1}\neq output_{2}\right)\leq 3b/\sqrt{2\pi s}$. Which translate to $(\varepsilon,\delta)=(0,\frac{3b}{\sqrt{2s\pi}})$ for differential privacy (DP), but remains an insufficient bound given that $\delta$ needs to be polynomial in inverse dataset size $1/s$ for DP[56].
Thm˜2 yields two immediate consequences: First, unless $m=\Omega(\sqrt{s})$, the retrofitting framework is resilient to poisoning attacks, and second in A/B testing case, privacy is guaranteed by design, covering prompts and outputs without additional mechanisms. These points are further discussed in App.˜F. It is important to note that, while other privacy preserving methods such as DP-SGD [46] enforce privacy by perturbing gradients and consequently suffer from utility degradation as $\varepsilon$ becomes small, BBoxER’s compression-based guarantees remain robust. These structural differences highlight a fundamental advantage of our approach.
@@ 5.1 Experiments
Experiment 4 (Robustness to data extraction). To assess the robustness of BBoxER against membership inference attacks (MIAs), we begin from the observation that most MIAs rely on the negative log-likelihood (NLL) of training samples as their primary signal to infer membership [65, 66, 4] - see also App.˜H for an exposition. Accordingly, we propose to measure the average of absolute differences in negative log-likelihood (NLL) between the base and the updated model for each training sample.
@@
We consider GSM8K as our training dataset, Llama-3.1-8B as our base model, and compute three metrics: Diff-full (NLL computed over the entire prompt), Diff-CoT-a (NLL computed over only the CoT and the answer), and the relative average exact match improvement on GSM8K-test from the base model. For a fair comparison, we evaluate BBoxER alongside standard fine-tuning and DP-AdamW, training all methods on GSM8K under different settings-varying budget, epochs, and parameter subsets (full details in App.˜H)—so that models are compared at similar utility. Our results, reported in Tab.˜2, demonstrate that at matched test accuracy, the NLL shifts under BBoxER are orders of magnitude smaller than under fine-tuning or DP-AdamW. This provides empirical confirmation of our theoretical claim in Sec.˜4.3: BBoxER is significantly more robust to extraction and membership inference attacks.
@@ 5.2 Discussion of experimental results
Computational cost. A limitation of our approach is that each iteration requires a full pass over the dataset to generate a comparison result. However, BBoxER remains memory-efficient, as it relies solely on LLM inference without the need to compute and store gradients, or large optimizer states. This allows us to fit a large batch size during training. The overhead introduced by the internal state of the BBO algorithm is negligible compared to the computational cost of LLM inference.
@@ Appendix C Theoretical Background: A Brief Overview of Statistical Learning Theory
C.1 Non Asymptotic Large Deviation Bounds from the 60s: Hoeffding, Bennett, Bernstein
@@
In case we have a prior on a bound $\sigma^{2}$ on the variance of a randomly drawn individual loss, we can use the bound of Bennett and Bernstein [104]:
@@ C.2.1 Bounds on the Generalization Loss based on Bonferroni
Thm˜1 extends this to more sophisticated cases, with $N(\omega,a,b)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\prod_{i=1}^{b}}k_{i}(\omega,a,b)>>b$ (in many cases, exponential in $b$ e.g. $N(\omega,a,b)\leq 2^{b}$ and $k_{i}\leq 2$) instead of $N(\omega,a,b)=b$. In spite of being larger, this bound, exponential in $b$, is tighter than bounds based on capacity.
@@ D.1 Evolutionary Algorithms
CMA-ES [13, 14, 15] iteratively updates a multivariate Gaussian random variable, initialized at the identity and parametrized by a covariance matrix and its variance (the so-called step-size). Each update is based on the correlations observed in successful mutations. This allows CMA-ES to adapt its step-size and search directions (the covariance matrix) to the problem landscape. Despite its $o(d^{2})$ computational cost and memory usage, due to the handling of the full covariance matrix, it is widely regarded as a powerful and reliable optimizer.
@@ E.2 Generalization Bounds for Randomized BBoxER
In contrast to random search, where all candidate parametrizations are chosen independently of their evaluation, comparison-based algorithms build candidates iteratively based on performance. If $\lambda$ candidates are evaluated at each step and the best is retained, then the number of internal states after one iteration is $\lambda$; after two iterations, $\lambda^{2}$; and after $n=b/\lambda$ iterations, at most $\lambda^{n}$. The total risk is, therefore, bounded by $\lambda^{n}\cdot\delta_{1,\epsilon}$, where $n$ is the number of iterations with $k_{i}>1$ (with $k_{i}=\lambda$ in such iterations); see [108].
@@ E.7 Dependence on Dataset Size: Scaling and Numerical Application
If we prefer the Bennett&Bernstein inequality (based on an assumption on the standard deviation, see Sec.˜C.1 [102, Section 8.2]), we get
@@
This equation, for $s=8000$ examples (close to GSM8k), $\delta=\frac{1}{2}$ (we are considering the median case), $\epsilon=0.01$ (we want that precision) and $\sigma=0.06$ (the standard variation of the loss is 0.06, we assume that the initial model is already good), leads to the claim that a budget $b\leq 91$ cannot lead to overfitting. If we assume $\epsilon=0.04$ and $\sigma=0.3$, we get with Hoeffding a bound $34$ and with Bennett&Bernstein a bound $88$. With $\epsilon=0.06$, $\sigma=0.3$, Bennett&Bernstein leads to a budget $189$. The budget becomes $482$ if we use $\epsilon=0.1$, $\sigma=0.3$. Also, these bounds scale linearly with $s$ (the number of user preferences in A/B testings or dataset size). These numbers are low, but we see that our bounds are not vacuous even in realistic scenarios.
@@ F.1 Privacy by Design
Algorithm stability.
(general case of applying BBoxER on a dataset $D$) We may also be interested in the influence of a single satisfaction response $r_{i}$ associated with a specific user, or in the case were we compute this preference signal from the data. In this setting, we consider the case where $D_{1}$ and $D_{2}$ differ by exactly one such response, altered during a run. Applying Equation˜7 with $m=1$, we get that $A(D_{1})$ and $A(D_{2})$ are the same with probability at least $1-3b/\sqrt{2s\pi}$ when we have $s$ data points and $b$ iterations. Which represents $\varepsilon=0$ and $\delta=3b/\sqrt{2s\pi}$ in Equation˜A.11 but is insufficient from a DP standpoint given that $\delta$ needs to be polynomial in $\frac{1}{s}$. A differentially private version of BBoxER can be obtained by adding a Gaussian noise to every user preference signal in case of A/B testing, or by adding a Gaussian noise to the decision metric (like accuracy) in case we compute the performance from the prompts/outputs of the model.